OpenStudy (anonymous):

How do I find the slope to 2x+3y=9

OpenStudy (anonymous):

\[2x+3y=9\]\[2x-9=-3y\]\[\frac{ -2 }{ 3 }x+3=y\] The equation for a line is y=mx+b m is the slope.

OpenStudy (anonymous):

so \[\frac{ -2 }{ 3 }\] is my slope?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

TY